In Fall 2013, we spotlighted 24 jackrabbits along 32-km of roads in grasslands in northern Colorado, measuring the perpendicular distance (in m) from the truck to each rabbit seen (data are shown in the table to right). Based on previous work, we know that we can reliably locate rabbits if they are within 45 m of either side of the road, then detectability drops off exponentially with increasing distance from the road. We are also confident that our linear spotlighting route effectively samples the rabbit population on the entire 62-km2 study area.

a) What is your naïve estimate of population density in 2013?

b) Estimate the size and density of the jackrabbit population in our study area in Fall 2013.

c) Based on two decades of annual counts, the jackrabbit population has fallen dramatically since its peak in Fall 2006, when we estimated that there were 1,902 rabbits in the study area.
At what annual rate has the rabbit population changed since 2006?

d) At this rate, in what year will the population crash, i.e. to less than 1 individual?
Answer #1

a. As we are seeing only 45 m to either side of road and travelled 32 kms and found 24 rabbits.

So, the total land area we searched for rabbit is = (32 km)*2*(45 m) = 2.88 km2.

Population density = no. of rabbits/ land area = 24/2.88 = 8.33 = 8 rabbits per km2.

b. Size in study area is = 8.33 * 62 = 516.67 = 517 rabbits

Density in our study area = 516.67/62 = 8.33 = 8 rabbits per km2same as in no. a

c. No. of years = 7 (2013-2006 = 7

Annual rate = (no. of rabbits in 2006 – no. of rabbits in 2013) / no. of years

                    = (1902 – 517)/7 = 197.85 = 198 rabbits per year decreased

    The rabbit population has decreased at 198 rabbits per year since 2006.

d. To have less than I individual let take no. of rabbits as 0.44<1

No of years required to have less than 1 individual = (517 – 0.44)/197.85 =2.61 years = 3 years approximately.

At this rate, in year 2016 will the population crash, i.e. to less than 1 individual.